Perdon Corporation manufactures safes—large mobile safes, and large walk-in stationary bank safes. As part of its annual budgeting process, Perdon is analyzing the profitability of its two products. Part of this analysis involves estimating the amount of overhead to be allocated to each product line. The information shown below relates to overhead.

Mobile Safes

Walk-in Safes

Units planned for production 200 60
Material moves per product line 310 190
Purchase orders per product line 450 350
Direct labor hours per product line 800 1,710

Part 1

New attempt is in progress. Some of the new entries may impact the last attempt grading.Your answer is partially correct.

The total estimated manufacturing overhead was $277,000. Under traditional costing (which assigns overhead on the basis of direct labor hours), what amount of manufacturing overhead costs are assigned to: (Round answers to 2 decimal places, e.g. 12.25.)

(1) One mobile safe

$enter a dollar amount per unit rounded to 2 decimal places

per unit
(2) One walk-in safe

$enter a dollar amount per unit rounded to 2 decimal places

per unit

eTextbook and Media

Part 2

Incorrect answer iconYour answer is incorrect.

The total estimated manufacturing overhead of $277,000 was comprised of $166,000 for materials handling costs and $111,000 for purchasing activity costs. Under activity-based costing (ABC): (Round answers to 2 decimal places, e.g. 12.25.)

What amount of materials handling costs are assigned to:

(a) One mobile safe

$enter a dollar amount per unit rounded to 2 decimal places

per unit
(b) One walk-in safe

$enter a dollar amount per unit rounded to 2 decimal places

per unit

eTextbook and Media

Part 3

Incorrect answer iconYour answer is incorrect.

The total estimated manufacturing overhead of $277,000 was comprised of $166,000 for materials handling costs and $111,000 for purchasing activity costs. Under activity-based costing (ABC): (Round answers to 2 decimal places, e.g. 12.25.)

What amount of purchasing activity costs are assigned to:

(a) One mobile safe

$enter a dollar amount per unit rounded to 2 decimal places

per unit
(b) One walk-in safe

$enter a dollar amount per unit rounded to 2 decimal places

per unit

eTextbook and Media

Part 4

Incorrect answer iconYour answer is incorrect.

Compare the amount of overhead allocated to one mobile safe and to one walk-in safe under the traditional costing approach versus under ABC. (Round answers to 2 decimal places, e.g. 12.25.)

Traditional Costing

Activity-Based Costing

Mobile safe

$enter a dollar amount rounded to 2 decimal places

$enter a dollar amount rounded to 2 decimal places

Walk-in safe

$enter a dollar amount rounded to 2 decimal places

$enter a dollar amount rounded to 2 decimal places
Answer #1

The total estimated manufacturing overhead was $277,000. Under traditional costing (which assigns overhead on the basis of direct labor hours), what amount of manufacturing overhead costs are assigned to: (Round answers to 2 decimal places, e.g. 12.25.)

(1) One mobile safe 277000*800/2510 = 88287 88287/200 = 441.43
(2) One walk-in safe 277000*1710/2510 = 188713 188713/60 = 3145.22

The total estimated manufacturing overhead of $277,000 was comprised of $166,000 for materials handling costs and $111,000 for purchasing activity costs. Under activity-based costing (ABC): (Round answers to 2 decimal places, e.g. 12.25.)

What amount of materials handling costs are assigned to:

(a) One mobile safe 166000/500*310 = 102920 102920/200 = 514.60
(b) One walk-in safe 166000/500*190 = 63080 63080/60 = 1051.33

The total estimated manufacturing overhead of $277,000 was comprised of $166,000 for materials handling costs and $111,000 for purchasing activity costs. Under activity-based costing (ABC): (Round answers to 2 decimal places, e.g. 12.25.)

What amount of purchasing activity costs are assigned to:

(a) One mobile safe 111000/800*450 = 62437.50 62437.50/200 = 312.19
(b) One walk-in safe 111000/800*350 = 48562.50 48562.50/60 = 809.38

Compare the amount of overhead allocated to one mobile safe and to one walk-in safe under the traditional costing approach versus under ABC. (Round answers to 2 decimal places, e.g. 12.25.)

Traditional Costing

Activity-Based Costing

Mobile safe 441.43 826.79
Walk-in safe 3145.22 1860.71